how do you solve for
ln(x+1) - ln(x-3)=ln(x)
thanks

Answer by user_dude2008(1861)

ln(x+1) - ln(x-3)=ln(x)

ln((x+1)/(x-3))=ln(x)


(x+1)/(x-3) = x

x+1 = x(x-3)

x+1 = x^2-3x

x^2-4x-1 = 0

x = 2 + sqrt(5) <-------------- Answer